OpenStudy (anonymous):

If there is an inverse relationship between x and y, and the constant of variation is forty, what is y when x equals five? two hundred forty-five thirty-five 8

OpenStudy (anonymous):

y = k/x y = 40/x y = 40/5 x = 8